a reaction does 27 kJ of work pushing up a piston. if the reaction simultaneously gains 23 kJ of heat was is deltaE gor the reaction?
Answer #1

work = ws - 27 kJ heat = q = 23KJ DE = ? DE = qtw DE = 23 - 27 DE = -4KJ